OpenStudy (anonymous):

the record low temperature in arizona is 1.8 times the record low temperature in rhode island.if the record low temperature in rhode island is 10 degrees F.what is the record low temperature in arizona . write and solve an equation

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@helppls123

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

@magbak

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

Let x be the lowest recorded temperature in AZ. It's 1.8 times the lowest rec. temp. in RI. What is the lowest rec. temp. in RI? . Then x=1.8*( lowest rec. temp. in RI)

OpenStudy (anonymous):

So the equation will be 1.8 *x

OpenStudy (mathmale):

Please see my previous comment / post.

OpenStudy (anonymous):

But I don't understand what will the equations be if the question says if the record low low temperature in R.I is 10 what I'm I supposed to do with the 10

OpenStudy (mathmale):

That 10 degrees is the record low temperature in the State of Rhode Island. Do you know where RI and AZ are located within the USA?

OpenStudy (mathmale):

Which one do you think is the warmer state?

OpenStudy (anonymous):

arizona

OpenStudy (anonymous):

But what will the equation be where is the 10 supposed to go

OpenStudy (mathmale):

The lowest recorded temperature in RI is 10 degrees. Is the lowest rec. temp. in AZ higher or lower than that?

OpenStudy (anonymous):

It's 1.8 times the record low temp is it lower

OpenStudy (mathmale):

It's 1.8 times the record low temp in RI. Why do you say "it is lower?"

OpenStudy (anonymous):

Oh sorry

OpenStudy (mathmale):

Look: If you multiply 10 degrees by 1.8, you get what? That result is the record low temp in AZ.

OpenStudy (anonymous):

It equals 18

OpenStudy (mathmale):

What equals 18? Not trying to give you a hard time, but all these numbers need labels for identification. 18 degrees is the ..... 10 degrees is the ...

OpenStudy (anonymous):

Sorry 18 is the 1.8 in Arizona time the 10 which is the temp for R.I

OpenStudy (mathmale):

That's better. 18 degrees is the record low for AZ and is 1.8 times 10 degrees, the record low for RI.

OpenStudy (anonymous):

So ur telling me that's my Equation

OpenStudy (mathmale):

I typed that out for you about 15 minutes ago: Let x be the lowest recorded temperature in AZ. It's 1.8 times the lowest rec. temp. in RI. What is the lowest rec. temp. in RI? . ==> Then x=1.8*( lowest rec. temp. in RI) <== You have already found x. What's its value?

OpenStudy (anonymous):

@DDCamp

OpenStudy (anonymous):

What will the equation be

OpenStudy (mathmale):

==> x=1.8*( lowest rec. temp. in RI) <== or x=1.8 (10 degrees)